OpenStudy (anonymous):

x divided by negative 4 is greater than or equal to 7

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

equal

OpenStudy (anonymous):

\[\frac{ x }{-4 }\ge 7,-x \ge 28,x \le-28\]

OpenStudy (anonymous):

x is less than or equal to negative twenty-eight x is greater than or equal to negative twenty-eight x is less than or equal to twenty-eight x is greater than or equal to twenty-eight this is the list of answers

OpenStudy (anonymous):

you can see the answer.

OpenStudy (anonymous):

ya ty